xx RegridSyncTest

RegridSyncTest

Christian D. Ott <cott@aei.mpg.de>

April 14 2006

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

7 Interfaces

General Implements: regridsynctest

Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11 myregridtestgf 10 myregridtestgf x01  compact110
11                 10  x01 description11This is the regrid test grid function
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels113
11                 10  x01variable type11REAL

11

8 Schedule This section lists all the variables which are assigned storage by thorn Carpet/RegridSyncTest. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11x
11myregridtestgf[3]11x

11x  11x

11

Scheduled Functions

CCTKxINITIAL x regridsynctestxsetup x do something at inital x

11x11Language:11fortran
11 x 11Sync:  11myregridtestgf
11 x 11Type: 11function
11 x 11Writes: 11regridsynctest::myregridtestgf(everywhere)

11

CCTKxPOSTREGRID x regridsynctestxdoxsomething x do something at postregrid and try to sync x

11x11Language:11fortran
11 x 11Sync:  11myregridtestgf
11 x 11Type: 11function
11 x 11Writes: 11regridsynctest::myregridtestgf(everywhere)

11

CCTKxEVOL x regridsynctestxevolve x do something at evol x

11x11Language:11fortran
11 x 11Sync:  11myregridtestgf
11 x 11Type: 11function
11 x 11Writes: 11regridsynctest::myregridtestgf(everywhere)

11

CCTKxPOSTREGRID x regridsynctestxsync x dummy routine for sync-ing

11x11Before: 11regridsynctestxoutput
11 x 11Language:11fortran
11 x 11Sync:  11myregridtestgf
11 x 11Type: 11function

11

CCTKxPOSTREGRID x regridsynctestxoutput x do something at postregrid x

11x11Before: 11regridsynctestxdoxsomething
11 x 11Language:11c
11 x 11Type: 11function